# Are Vision Language Models Texture or Shape Biased and Can We Steer Them?

### v1

@@ Appendix F Ablation of Temperature Scaling
Appendix G Results on CLIP Models
In this section, we provide results for CLIP models that we referenced in the main paper. We provide results for different architectures under three different prompting strategies: a computation of zero-shot centroids from 80 different prompts including usage of the class name [11], "a photo of {class}." which is often used as a default prompt (note the dot), and "{class}" (without dot). We will argue that the latter is more comparable to the VQA task of our VLMs - but of course, VLMs may have a better representation in weights. Either way, the shape bias does not significantly deviate between the three strategies. Tab. 6 shows the obtained shape bias (and accuracy) measurements.
We also noticed that the observed scaling laws in [29] do not always hold for vision encoders, despite an increase in parameters from EVA02-CLIP-E/14+ (5B) to EVA02-CLIP-8B, we actually see a significant decrease in shape bias (but an improvement in accuracy).
@@ Appendix G Results on CLIP Models
EVA01-CLIP-g/14 [17] 80 Prompts [11] $66.03$ $87.83$
EVA01-CLIP-g/14 [17] "a photo of a {class}." $66.03$ $87.08$
EVA01-CLIP-g/14 [17] "{class}" $66.44$ $86.67$
EVA02-CLIP-8B@448px [82] 80 Prompts [11] $58.26$ $91.83$
EVA02-CLIP-8B@448px [82] "a photo of a {class}." $57.58$ $89.00$
EVA02-CLIP-8B@448px [82] "{class}" $56.60$ $88.33$
EVA02-CLIP-E/14+ [17] 80 Prompts [11] $65.62$ $90.67$
EVA02-CLIP-E/14+ [17] "a photo of a {class}." $64.44$ $89.75$
EVA02-CLIP-E/14+ [17] "{class}" $62.48$ $86.17$
CLIP-ViT-L/14 [11] 80 Prompts [11] $60.95$ $84.08$
CLIP-ViT-L/14 [11] "a photo of a {class}." $60.20$ $84.17$
CLIP-ViT-L/14 [11] "{class}" $60.16$ $81.17$
CLIP-ViT-L/14@336px [11] 80 Prompts [11] $61.52$ $86.83$
CLIP-ViT-L/14@336px [11] "a photo of a {class}." $60.56$ $86.42$
CLIP-ViT-L/14@336px [11] "{class}" $59.80$ $83.75$
CLIP-ResNet-50 [11] 80 Prompts [11] $19.70$ $77.83$
CLIP-ResNet-50 [11] "a photo of a {class}." $20.96$ $72.75$
CLIP-ResNet-50 [11] "{class}" $20.77$ $71.83$
CLIP-ResNet-101 [11] 80 Prompts [11] $25.50$ $74.83$
CLIP-ResNet-101 [11] "a photo of a {class}." $25.23$ $71.00$
CLIP-ResNet-101 [11] "{class}" $25.41$ $70.83$
Tab. 7 shows the delta between shape bias and accuracy for some selected VLMs and their vision encoders under zero-shot classification (using "{class}" as prompt). All selected models rely on frozen vision encoder during training. As discussed in Sec. 4.3, the VLM accuracy decreases in all cases, but shape bias can deviate in either direction.
Table 7: Quantitative comparison between shape bias and accuracy of VLMs (VQA) and their respective vision encoders (CLIP zero-shot classification) on cue-conflict.
VLM - Encoder [%]
VLM Vision Encoder Accuracy Shape-Bias
LLaVA v1.5 7B [61]  CLIP ViT-L/14@336px[11] $-3.50$ $4.2$
LLaVA v1.5 13B [61] $-3.00$ $1.5$
LLaVA-NeXT 7B [62] $-9.92$ $-3.9$
LLaVA-NeXT 13B [62] $-0.33$ $-2.7$
LLaVA-NeXT 34B [62] $-1.08$ $-0.2$
MoE-LLaVA v1.5 Phi2 x4 [67] $-1.42$ $-0.3$
MoE-LLaVA v1.5 Qwen x4 [67] $-24.25$ $3.0$
MoE-LLaVA v1.5 StableLM x4 [67] $-3.67$ $-0.8$
InstructBLIP FLAN-T5-XL [23]  EVA-01-CLIP ViT-g/14@224px [17] $-6.83$ $1.8$
InstructBLIP Vicuna-7B [23] $-14.42$ $7.4$
Emu2-Chat [66]  EVA-02-CLIP-E/14+@448px [17] $-11.08$ $-9.5$
Appendix H Results on ImageNet-trained Models
Complementary to the results on CLIP, we also provide some shape bias evaluations of ImageNet-trained/finetuned models in Tab. 8. Note how ViTs are much more shape-biased than ResNets, as shown in [29], and yet, after ImageNet-finetuning, the previously well-performing CLIP (ViT-L/14@336px) model drops from 59.80 % to just 32.1 % of shape bias.
@@ Appendix H Results on ImageNet-trained Models
ResNet-50 [59] 22.3 67.33
ResNet-50 (timm) [83] 23.1 65.42
ResNet-152 (timm) [83] 28.6 65.83
ViT-B/16 (ImageNet-21k pretraining) [32] 45.4 63.67
ImageNet-finetuned CLIP (ViT-L/14@336px) [83] 32.1 82.75
Appendix I Additional Thoughts on the Image Captioning Task
@@
While we assume that most embedding models will provide similar classification performance if the description clearly mentions one class, it is unclear how the classification is biased if the description refers to multiple classes, invalid classes, or is generic. Thus, we additionally ablate results with SFR-Embedding [84] which at the time of writing was the overall SOTA English embedding model on the Massive Text Embedding Benchmark (MTEB) [85]. While the accuracy improved by a negligible amount, shape bias results were largely unaffected. Thus, we settled for the faster ember-v1 models.
@@
Based on our LLM analysis, we notice that in the majority of cases (min: $79.3\%$, mean: $92.2\%$, median: $93.0\%$, max: $97.6\%$; minimum is given by InstructBLIP Flan-T5-xl [23]), descriptions do not refer to multiple labels, thus a potential bias of the embedding model is negligible for our analysis.
@@
In the cases where the caption is indeed generic, our choice of embedding model may have biased our study. However, the previously mentioned embedding with SFR-Embedding [84] showed similar trends. Thus, we assume that most other SOTA embedding models would behave similarly - yet, we are excited how future embedding models will embed these cases.
Another option is to remove generic responses from the analysis. We observe that this typically increases shape bias (and accuracy) - naturally more notably in cases where the generic ratio was high (Tab. 9). E.g., for the extreme case of GPT-4V [12], shape bias increases by $9.3\%$ and accuracy by $40.47\%$ (!). This preprocessing also seems to restore scaling laws to a large extent: larger models achieve higher shape bias and accuracy. One outlier to this trend is InternVL Chat v1.2+ [2]. It may be intriguing to replace the reported results in Sec. 4/Fig. 2 with the analysis on non-generic responses, but we avoid doing so, as this would a) remove a significant portion of results; b) lead to poorly comparable results obtained on different subsets.
Table 9: Comparison of shape bias and accuracy for the Image Captioning tasks for all responses and only responses which an LLM did not classify as generic.
@@
Appendix J Class-wise Texture/Shape Bias
Following the original shape bias study [1], we include plots that show the class-specific texture/shape bias for all our models in VQA and Image Captioning (Fig. 9).
(a) VQA
(b) Image Captioning
@@
Appendix K Responsibility to Human Subjects
We did not collect any human data ourselves. Instead, the human shape bias data is taken from [1] who collected the data in a controlled psychophysical laboratory (Wichmann-lab in Tübingen, Germany). Participants provided written informed consent and were compensated above minimum wage. The data was open-sourced by the authors on github without personally identifiable information. The experiment was conducted in accordance with institutional guidelines for human subjects research. For further information on the human experiment, we refer the reader to [1].
Appendix L Visualization of Vision Steering
We visualize the effects of vision steering (Sec. 5.1 in the main paper) in Fig. 10.

### v2

@@ Appendix H Ablation of Temperature Scaling
Appendix I Agreement sets on texture/shape cue-conflict
In this section, we aim to identify "agreement sets" of VLMs predictions on the texture/shape cue-conflict dataset (under the default VQA), i.e., we want to find samples where all 19 models behave the same to see if we can identify some common patterns. Specifically, we search for samples where all models predict the shape label, texture label, or generally make a correct or wrong prediction, respectively. For each agreement set, we show 4 random examples in Fig. 11. For this analysis alone, it is not evident to us when models predict texture or shape labels, respectively. Overall, the very low number of samples where all models make a false prediction (6 out of 1,200) is a good indication of the quality of the texture/shape cue-conflict dataset (Geirhos et al., 2019). We find that these error samples are still recognizable (but we may be biased, knowing the ground truth). There seems to be also a quite large subset of samples where all models make correct predictions (320 out of 1,200).
Refer to caption Refer to caption Refer to caption Refer to caption
Shape: bird clock bicycle airplane
Texture: cat bicycle dog bird
((a)) All models agree on shape (total $N=161/1200$).
Refer to caption Refer to caption Refer to caption Refer to caption
Shape: bird airplane keyboard boat
Texture: car chair clock car
((b)) All models agree on texture (total $N=38/1200$).
Refer to caption Refer to caption Refer to caption Refer to caption
Shape: bird bicycle airplane car
Texture: bear dot chair boat
((c)) All models predict correctly (total $N=320/1200$).
Refer to caption Refer to caption Refer to caption Refer to caption
Shape: bear keyboard bear knife
Texture: oven truck oven oven
((d)) All models predict falsely (total $N=6/1200$).
Figure 11: Agreement sets. We show 4 random samples where all VLMs show the same behavior under the default VQA prompt.
Appendix J Results on CLIP Models
In this section, we provide results for CLIP models that we referenced in the main paper. We provide results for different architectures under three different prompting strategies: a computation of zero-shot centroids from 80 different prompts including usage of the class name Radford et al. (2021), "a photo of {class}." which is often used as a default prompt (note the dot), and "{class}" (without dot). We will argue that the latter is more comparable to the VQA task of our VLMs - but of course, VLMs may have a better representation in weights. Either way, the shape bias does not significantly deviate between the three strategies. Tab. 7 shows the obtained shape bias (and accuracy) measurements.
We also noticed that the observed scaling laws in Geirhos et al. (2021) do not always hold for vision encoders, despite an increase in parameters from EVA02-CLIP-E/14+ (5B) to EVA02-CLIP-8B, we actually see a significant decrease in shape bias (but an improvement in accuracy).
@@ Appendix I Agreement sets on texture/shape cue-conflict
EVA01-CLIP-g/14 Sun et al. (2023b) 80 Prompts Radford et al. (2021) $66.03$ $87.83$
EVA01-CLIP-g/14 Sun et al. (2023b) "a photo of a {class}." $66.03$ $87.08$
EVA01-CLIP-g/14 Sun et al. (2023b) "{class}" $66.44$ $86.67$
EVA02-CLIP-8B@448px Sun et al. (2024) 80 Prompts Radford et al. (2021) $58.26$ $91.83$
EVA02-CLIP-8B@448px Sun et al. (2024) "a photo of a {class}." $57.58$ $89.00$
EVA02-CLIP-8B@448px Sun et al. (2024) "{class}" $56.60$ $88.33$
EVA02-CLIP-E/14+ Sun et al. (2023b) 80 Prompts Radford et al. (2021) $65.62$ $90.67$
EVA02-CLIP-E/14+ Sun et al. (2023b) "a photo of a {class}." $64.44$ $89.75$
EVA02-CLIP-E/14+ Sun et al. (2023b) "{class}" $62.48$ $86.17$
CLIP-ViT-L/14 Radford et al. (2021) 80 Prompts Radford et al. (2021) $60.95$ $84.08$
CLIP-ViT-L/14 Radford et al. (2021) "a photo of a {class}." $60.20$ $84.17$
CLIP-ViT-L/14 Radford et al. (2021) "{class}" $60.16$ $81.17$
CLIP-ViT-L/14@336px Radford et al. (2021) 80 Prompts Radford et al. (2021) $61.52$ $86.83$
CLIP-ViT-L/14@336px Radford et al. (2021) "a photo of a {class}." $60.56$ $86.42$
CLIP-ViT-L/14@336px Radford et al. (2021) "{class}" $59.80$ $83.75$
CLIP-ResNet-50 Radford et al. (2021) 80 Prompts Radford et al. (2021) $19.70$ $77.83$
CLIP-ResNet-50 Radford et al. (2021) "a photo of a {class}." $20.96$ $72.75$
CLIP-ResNet-50 Radford et al. (2021) "{class}" $20.77$ $71.83$
CLIP-ResNet-101 Radford et al. (2021) 80 Prompts Radford et al. (2021) $25.50$ $74.83$
CLIP-ResNet-101 Radford et al. (2021) "a photo of a {class}." $25.23$ $71.00$
CLIP-ResNet-101 Radford et al. (2021) "{class}" $25.41$ $70.83$
Appendix K Results on ImageNet-trained Models
Complementary to the results on CLIP, we also provide some shape bias evaluations of ImageNet-trained/finetuned models in Tab. 8. Note how ViTs are much more shape-biased than ResNets, as shown in Geirhos et al. (2021), and yet, after ImageNet-finetuning, the previously well-performing CLIP (ViT-L/14@336px) model drops from 59.80 % to just 32.1 % of shape bias.
@@
ResNet-50 He et al. (2015) 22.3 67.33
ResNet-50 (timm) Wightman (2019) 23.1 65.42
ResNet-152 (timm) Wightman (2019) 28.6 65.83
ViT-B/16 (ImageNet-21k pretraining) Dosovitskiy et al. (2021) 45.4 63.67
ImageNet-finetuned CLIP (ViT-L/14@336px) Wightman (2019) 32.1 82.75
Appendix L Additional Thoughts on the Image Captioning Task
@@
While we assume that most embedding models will provide similar classification performance if the description clearly mentions one class, it is unclear how the classification is biased if the description refers to multiple classes, invalid classes, or is generic. Thus, we additionally ablate results with SFR-Embedding Meng et al. (2024) which at the time of writing was the overall SOTA English embedding model on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023). While the accuracy improved by a negligible amount, shape bias results were largely unaffected. Thus, we settled for the faster ember-v1 models.
@@
Based on our LLM analysis, we notice that in the majority of cases (min: $79.3\%$, mean: $92.2\%$, median: $93.0\%$, max: $97.6\%$; minimum is given by InstructBLIP Flan-T5-xl Dai et al. (2023)), descriptions do not refer to multiple labels, thus a potential bias of the embedding model is negligible for our analysis.
@@
In the cases where the caption is indeed generic, our choice of embedding model may have biased our study. However, the previously mentioned embedding with SFR-Embedding Meng et al. (2024) showed similar trends. Thus, we assume that most other SOTA embedding models would behave similarly - yet, we are excited how future embedding models will embed these cases.
Another option is to remove generic responses from the analysis. We observe that this typically increases shape bias (and accuracy) - naturally more notably in cases where the generic ratio was high (Tab. 9). E.g., for the extreme case of GPT-4V OpenAI (2023), shape bias increases by $9.3\%$ and accuracy by $40.47\%$ (!). This preprocessing also seems to restore scaling laws to a large extent: larger models achieve higher shape bias and accuracy. One outlier to this trend is InternVL Chat v1.2+ Chen et al. (2024). It may be intriguing to replace the reported results in Sec. 4/Fig. 2 with the analysis on non-generic responses, but we avoid doing so, as this would a) remove a significant portion of results; b) lead to poorly comparable results obtained on different subsets.
Table 9: Comparison of shape bias and accuracy for the Image Captioning task for all responses and only responses which an LLM did not classify as generic.
@@
Appendix M Ablation of multi-modal training stages
Table 10: Comparison of LLaVA v1.5-7B models between Stage 1 and Stage 2 with additional non-generic metrics.
All responses Non-generic
Shape Accur- Avg. Single Class Generic Shape Accur-
Model Bias [%] acy [%] Tokens Ratio [%] Ratio [%] Bias [%] acy [%]
Stage 1 61.8 73.25 143.5 54.1 31.4 64.5 90.32
Stage 2 61.4 76.08 12.1 73.8 19.2 62.8 87.24
LLaVA models are trained in two stages. During Stage 1 training the vision encoder and LLM remain frozen and only the parameters of the connector in between are updated. During Stage 2, the parameters of the LLM are included as well. To ablate the effect of these different training stages on our results, we repeated our experiments with a Stage 1 LLaVA-1.5-Vicuna-7B checkpoint in comparison to the final Stage 2 model and show the results in Tab. 10.
It is worth noting, that due to the lack of proper instruction-tuning, the Stage 1 model does not follow the VQA instructions and depending on the prompt either generates gibberish (e.g., ‘‘The image is a collage of various items, including a bottle, a jar, a can, a spoon, a fork, a knife, [..keeps repeating..]’’) or is consistently giving a wrong prediction. However, we can assess the bias in the captioning setting. These answers are repetitive and noisy, too, but can still be discriminated by the embedding models.
We observe the following trends: instruction tuning (Stage 2) reduces the verbosity of generated descriptions (avg. tokens) and increases accuracy. The shape bias of all responses is only marginally affected (slightly decreases). The instruction-tuned model generates significantly fewer generic captions (i.e., those that do not refer to any label), but it also reduces the ratio of answers referring to multiple classes and, thus, becomes more biased by forgetting one cue. Because we force the sentence embedding models to make a prediction even on generic captions, we may bias the results. To compensate for this we repeated the analysis only on the non-generic captions (similar to Tab. 9).
For non-generic responses the Stage 1 model performs better (higher accuracy) and achieves a higher shape bias. Taken together, this indicates that instruction tuning (at least in this specific LLaVA model) seems to force the model to make stronger predictions (i.e., more correct predictions but also forgetting one cue) and increases texture bias.
Appendix N Does multi-modal training guarantee shape-biased encoders?
We aim to understand a critical aspect of encoder training: the combination of vision and language. It was already demonstrated that ViT-CLIP models show an increased shape bias in comparison to vision-only models independent of their architecture, training data, or method Geirhos et al. (2021). This may suggest that joint embedding alone increases shape bias. To verify this hypothesis, we measure the shape bias of a ResNet-50-based CLIP Radford et al. (2021) - and observe an opposite trend. With just $20.8\%$ (refer to Appendix J for details), this model even slightly reduced shape bias compared to an ImageNet-trained ResNet-50 ($22.2\%$). As such, just fusing language into encoder training does not guarantee an increased shape bias and the results strongly depend on the vision architecture, as well. For the design of shape-biased VLMs, it is, thus, reasonable to rely on representations of ViT-based CLIP as opposed to vision-only models.
Appendix O Other biases in deep neural networks
We want to emphasize that we deliberately exclude high-level, societal biases from our considerations. Our study merely considers biases in the sense of low-level feature-based cues preferences.
High-level vision biases have been widely investigated, such as single-demographic effects (race and gender) for face recognition tasks Buolamwini & Gebru (2018); Raji & Buolamwini (2019). For language models, several works focus on investigating societal biases, such as gender and race Barikeri et al. (2021); Lauscher et al. (2021) and ways of debiasing them Lauscher et al. (2021); Meade et al. (2022); Guo et al. (2022), or explicitly forcing them Haller et al. (2023). A recent study also found that LLMs are biased towards high-value over likely options Sivaprasad et al. (2024). Another study focused on encoded moral beliefs Scherrer et al. (2023). LLMs can also pick up human traits - one study found that adding "take a deep breath" to prompts improves performance Yang et al. (2024). Of course, some of the uni-modal biases also apply to VLMs (e.g., Yang et al. (2024)), but a few works have also explicitly focused on biases in VLMs. For example, neurons of CLIP Radford et al. (2021) were studied in Goh et al. (2021), revealing that some neurons respond to the same concept regardless of its presentation, which is a potential reason for the high generalizability. On the other hand, this enables attacks by rendering text on images (typographic attacks). Additionally, several works demonstrated that VLMs fail to count objects Radford et al. (2021); Liu et al. (2021); Thrush et al. (2022), and generally struggle in structured tasks Zhai et al. (2022a).
Appendix P Class-wise Texture/Shape Bias
Following the original shape bias study Geirhos et al. (2019), we include plots that show the class-specific texture/shape bias for all our models in VQA and Image Captioning (Fig. 12).
((a)) VQA
((b)) Image Captioning
@@
Appendix Q Error Consistency Plot
Figure 13 shows the error consistency between our predictions of our VLMs in the VQA, CLIP models including their encoders, ImageNet-trained or fine-tuned models, and ten human subjects.
Figure 13: Error consistency between different observers on texture/shape cue-conflict. VLMs make similar errors on the cue-conflict dataset and share similarities with their vision encoders. In terms of errors, VLMs are also more similar to humans than ImageNet-trained/finetuned models. We measure the pair-wise error consistency Geirhos et al. (2020b) between predictions. For this analysis, an error is any answer that does not belong to the shape class (analogous to Geirhos et al. (2021)). Shown responses belong to LLM-based VLMs (under the VQA task), other selected models including ImageNet models, (some) VLM encoders under ImageNet-finetuning and zero-shot classification, and ten human subjects.
Appendix R Responsibility to Human Subjects
We did not collect any human data ourselves. Instead, the human shape bias data is taken from Geirhos et al. (2019) who collected the data in a controlled psychophysical laboratory (Wichmann-lab in Tübingen, Germany). Participants provided written informed consent and were compensated above minimum wage. The data was open-sourced by the authors on GitHub without personally identifiable information. The experiment was conducted in accordance with institutional guidelines for human subjects research. For further information on the human experiment, we refer the reader to Geirhos et al. (2019).
Table 11: Example conversation for LLM-guided Prompt Search. Mock conversation.
User Your task is to help me design an ideal image classification prompt for a vision/language model to MINIMIZE the shape bias without significantly hurting accuracy. The provided photos show natural objects modified to contain shape and texture of conflicting object classes. E.g., it could be a photo of an elephant (shape) with dog fur (texture). The model should classify the image as ’dog’ based on the texture of the object and ignore the shape. The model’s accuracy is the percentage of correctly classified images. The shape bias is the ratio of how often the model classified based on shape over texture. You can test your prompt by outputting a single new line starting with ’PROMPT: ’. Do not list options - the system will provide them automatically. Try to keep the prompt as short and simple as possible but be creative. It might be reasonable to summarize insights of previous attempts and to outline your goals before responding with a new prompt, but make sure that only the prompt starts with ’PROMPT:’. In response to the prompt you will be told the accuracy and shape bias. Then you will refine the prompt and we will continue until I say stop. Let’s go!
Assistant PROMPT: Which option best describes the image?
User Prompt: […], Accuracy: 82.58 %, Shape Bias: 59.43 %. What is your next prompt?
Assistant To minimize the shape bias while maintaining accuracy, let’s encourage the model to focus more on textures than shapes. Here’s my next prompt: PROMPT: Based on the surface characteristics and patterns, which category does the image most resemble?
User Prompt: […], Accuracy: 78.25 %, Shape Bias: 50.9 %. Can you improve this?
$\cdots$
Assistant I have provided various prompts and explored different approaches to minimize the shape bias while maintaining high accuracy. The current optimal prompt is: […] Here are some additional suggestions for minimizing the shape bias: […] These methods may require more computational resources and time investment. If you have other questions or need guidance on different topics, please don’t hesitate to reach out. Wishing you the best in your continued efforts! At this point, I believe that we have explored various approaches, and the current prompt is the best-performing one among those tested. If you decide to pursue any of the mentioned alternatives, I am confident that you will find ways to improve the model’s performance and minimize the shape bias. Good luck with your continued work! As our conversation has concluded, I would like to thank you for engaging in this exercise, and I hope I have been helpful in providing prompts and suggestions. If you have other questions or need guidance on different topics in the future, please don’t hesitate to reach out. Wishing you the best in your projects and research!